Solve each system by the substitution method.
x=5y-3
x=8y+4

Answer by Alegend(2)

x = 5y - 3
x = 8y + 4
Rewrite the first equation:
x + 3 = 5y
y = (x + 3)/5
Substitute this into the second equation:
x = 8((x + 3)/5) + 4
Simplify:
x = (8x + 24)/40 + 4
Solve:
5.75 = x
And there ya go! ;)